# REAP the Experts: Why Pruning Prevails for One-Shot MoE compression

### v1

@@ 6 Discussion
Similar to prior work, we find that expert merging performs reasonably well on MC benchmarks. This may be because MC tasks only require a discriminative function that can be approximated by an average expert. In contrast, merging fails to maintain model quality on generative tasks, particularly at 50% compression. Generative tasks require auto-regressive generation, a capability that is lost when the router’s fine-grained control is removed. Compared to expert pruning, merging is less consistent, exhibiting higher variance across models and compression ratios. The outputs of expert merged models are more repetitive and less closely aligned with the base model compared with pruned models. Taken together, these observations are direct evidence of alterations to the functional manifold of the SMoE layers discussed in Section˜3.3 stemming from the loss of the router’s input-dependent control over the experts and subsequent introduction of novel functions due to tying of the merged expert gates.
Overall, expert pruned models offer consistently higher accuracy than merged models on generative tasks. REAP is a robust pruning criterion that generalizes across a wide array of SMoE architectures, compression ratios, and generative tasks. By taking into consideration both the router gate-values and expert activation norms, REAP prunes the experts which contribute the least to each layers output on a per-token average, regardless of usage frequency. REAP is scalable, achieving near-lossless compression on coding tasks with Qwen3-Coder-480B and Kimi-K2. The successes of REAP highlights the crucial importance of preserving coordination between the router and experts. Compression methods which impair the router’s ability to independently modulate expert outputs or tie gate-values are less likely to succeed.
@@ 7 Conclusion
Our analysis of current SMoE expert merging techniques finds that the router’s loss of independent control over experts results in functional subspace collapse. In contrast, expert pruning produces a coordinate subspace of the original layer which maintains the topology of the functional manifold. Based on our findings that the coordination between the router and experts is fundamental, we introduce REAP, a novel expert pruning method which prunes experts that contribute the least to the layer’s output. Empirically, we demonstrate that REAP retains remarkably high accuracy on an wide array of generative tasks across a diverse set of model architectures. We hope that this work inspires further compression techniques for SMoEs and facilitates the deployment of accurate, domain-specific models in resource constrained settings.
@@ Acknowledgments
We would like to acknowledge the helpful feedback of Mohammed Adnan and Rohan Jain. ML and YI gratefully acknowledge the support of Alberta Innovates (ALLRP-577350-22, ALLRP-222301502), the Natural Sciences and Engineering Research Council of Canada (NSERC) (RGPIN-2022-03120, DGECR-2022-00358), and Defence Research and Development Canada (DGDND-2022-03120). ML and YI are grateful for computational resources made available to us by the Digital Research Alliance of Canada. YI is supported by a Schulich Research Chair.
@@
This work research focused on the algorithmic compression of SMoE models and does not involve the use of human subjects, personally identifiable information, or sensitive data. The datasets used for calibration and evaluation (e.g., c4, evol-codealpaca) are publicly available. Our aim is enable the use of large-scale SMoE models in resource constrained settings. However, we acknowledge that compression techniques such as REAP could potentially facilitate deployment of models for malicious purposes. Further, our compression methods are applied to pre-trained models and any biases related to fairness, discrimination, or representation inherent in the original models may be present in their compressed versions. We make no attempt in this work to mitigate these potential biases. The primary contribution of this paper is technical, and we do not foresee any new, direct ethical concerns arising from our proposed methodology beyond those already associated with the deployment of large language models.
@@ Appendix A Extension to Hierarchical Clustering
While Theorem 1 analyses pairwise merging, practical implementations often employ hierarchical clustering to form groups of experts. Consider a cluster $C=\{f_{i_{1}},\ldots,f_{i_{k}}\}$ of $k$ experts merged into a single representative $\tilde{f}_{C}$. The original contribution of this cluster can be decomposed as:
@@
After hierarchical merging, the router must apply the summed gate $\sum_{j\in C}g_{i_{j}}$ to a single, static cluster representative $\tilde{f}_{C}$, typically computed as a weighted average of the cluster members based on calibration data. This induces an irreducible error:
Theorem 2 (Hierarchical clustering error).
@@ Appendix B Additional empirical evidence for loss of independent control
In Figure˜1(a), Qwen3’s layer 0 exemplifies the contraction of the functional output space by merging in early layers. The original 128 experts span from 0.4-0.4 to $1.0$ along PC1, pruning maintains this full range with 64 experts, while merging contracts the distribution to approximately [0.2,0.3][-0.2,0.3], a 5-fold reduction. This contraction is dramatic in late layers, where experts are more specialized. As depicted in Figure˜4(f), the original 15 experts of Llama-4’s layer 47 occupy a vast, multi-modal space spanning PC1 coordinates from 800-800 to $600$. Pruning preserves this remarkable diversity, with the 8 surviving experts distributed across the same multi-modal regions. However, merging induces a catastrophic collapse to a tiny cluster around coordinates $(200,0)$, representing nearly two orders of magnitude reduction in functional diversity. This pattern intensifies with the number of experts: Qwen3’s layer 47 (Figure˜1(b)) shows the most severe collapse, with 128 original experts spanning PC1 from 200-200 to $300$ reduced to a minute region after merging, while its 64 pruned experts maintain the original distribution’s full breadth.
Manifold geometry preservation
Across all models and layers, we observe a fundamental geometric principle: pruning preserves the topology of the functional manifold while merging fundamentally alters it. This distinction is most clearly visible in ERNIE’s representations (Figures 4(a) and 4(b)). In layer 1, the original 64 routed experts plus 2 shared experts form a characteristic curved structure with several outliers representing specialized experts. After pruning, the red points precisely overlay the gray ghost of the original distribution, including the outlier positions, demonstrating that each surviving expert maintains its exact functional role. The merged configuration, however, shows all experts collapsed into a tight cluster at the distribution’s centre, eliminating both the outliers and the manifold’s curvature.
The preservation of manifold geometry under pruning reflects the mathematical structure of the operation: the pruned hypothesis class is a coordinate subspace of the original, with the router maintaining independent control over each surviving expert. The geometric collapse under merging visualizes the loss of independent control when gates $g_{i}$ and $g_{j}$ are tied into their sum $(g_{i}+g_{j})$, the router can no longer independently modulate the two underlying functions, forcing the model to approximate the dynamic mixture $r(x)f_{i}(x)+(1-r(x))f_{j}(x)$ with a static expert $\tilde{f}_{\alpha}$.
Mixtral, with only 8 experts, provides an interesting edge case (Figures 4(c) and 4(d)). Even with fewer experts, the same geometric principles apply. Pruning maintains the convex hull of the original distribution while merging contracts it. The less dramatic collapse compared to models with more experts suggests that with fewer experts, each must remain more general, leading to lower $\|\Delta_{ij}\|^{2}$ (expert gap) and lower $\mathrm{Var}[r(x)]$ (policy variability), both factors in our irreducible error bound.
@@
(e) Llama-4 Layer 0
(f) Llama-4 Layer 47
Figure A4: (4(a),4(c),4(e)) Functional subspace (PCA) for early SMoE layers. Pruning (blue) preserves the manifold geometry; merging (green) collapses it toward the centre. (4(b),4(d),4(f)) Functional subspace (PCA) for late MoE layers.
@@ Appendix C Evaluation details
Mathematical reasoning is assessed on GSM8K (Cobbe et al., 2021) and MATH-500 (Hendrycks et al., 2021b; Lightman et al., 2023) benchmarks using the evalscope (ModelScope Team, 2024) framework. To assess creative writing, we use 146 creative writing prompts sampled from WildBench (Lin et al., 2024) with GPT-4o used as the judge to evaluate the model responses. We report normalized scores using the WildBench rubric.
@@ Appendix D Baseline methods
The clusters are merged with equation 14.
Appendix E Additional results
Table˜A4 shows the full suite of MC question answering benchmarks and the average result across all models and methods. Table˜A5 tabulates code generation accuracy of compressed SMoE models calibrated on evol-codealpaca. Eval+ is the average of MBPP, MBPP+, HumanEval (HE), HE+. The Code Avg column is the average of Eval+ and LiveCodeBench (LiveCode). Table˜A6 summarizes the accuracy of the various compression methods studied when calibrated with the c4 dataset on coding and MC benchmarks. Notably, while the MC performance is generally slightly higher than models calibrated on evol-codealpaca, the resulting code generation quality is abysmal, with most models failing to generate coherent output.
@@ Appendix E Additional results
REAP 0.344 $\pm$ 0.004 0.504 $\pm$ 0.008 0.745 $\pm$ 0.005 0.489 $\pm$ 0.013 0.507 $\pm$ 0.005 0.311 $\pm$ 0.003 0.625 $\pm$ 0.031 0.623 $\pm$ 0.007 0.518 $\pm$ 0.004
Mixtral-8x7B-
Instruct-v0.1
Baseline 0.650 0.842 0.887 0.861 0.691 0.496 0.722 0.740 0.736
@@
REAP 0.544 $\pm$ 0.005 0.785 $\pm$ 0.005 0.837 $\pm$ 0.003 0.778 $\pm$ 0.002 0.554 $\pm$ 0.001 0.462 $\pm$ 0.005 0.715 $\pm$ 0.013 0.679 $\pm$ 0.005 0.669 $\pm$ 0.001
Llama-4-Scout-
17B-16E-
Instruct
Baseline 0.627 0.848 0.879 0.823 0.803 0.462 0.765 0.692 0.738
@@
REAP 0.427 0.604 0.662 0.642 0.569 0.318 0.606 0.640 0.559
Qwen3-Coder-
480B-A35B-
Instruct-FP8
Baseline 0.644 0.822 0.906 0.841 0.850 0.468 0.751 0.717 0.750
@@
Kimi-K2-
Instruct-
W4A16
Baseline 0.712 0.879 0.913 0.765 0.872 0.504 0.783 0.811 0.780
@@
Table A5: Detailed benchmark results for non-agentic code generation tasks. Eval+ is the average of MBPP, MBPP+, HE, HE+. The Code Avg column is the average of Eval+ and LiveCodeBench (LiveCode).

### v2

@@ 6 Discussion
Similar to prior work, we find that expert merging performs reasonably well on MC benchmarks. This may be because MC tasks only require a discriminative function that can be approximated by an average expert. In contrast, merging fails to maintain model quality on generative tasks, particularly at 50% compression and high-granularity architectures. Generative tasks require auto-regressive generation, a capability that is impaired when the router’s fine-grained control is removed or novel expert functions are introduced. Compared to expert pruning, merging is less consistent, exhibiting higher variance across models and compression ratios. The outputs of expert merged models are more repetitive and less closely aligned with the base model compared with pruned model’s outputs. Taken together, these observations are direct evidence of functional manifold distortion of the SMoE layers discussed in Section˜3.3.
Overall, expert pruned models offer consistently higher accuracy than merged models on generative tasks. REAP is a robust pruning criterion that generalizes across a wide array of SMoE architectures, compression ratios, and generative tasks. By taking into consideration both the router gate-values and expert activation norms, REAP minimizes the reconstruction error bound by pruning experts which contribute the least to each layers output. REAP is scalable, achieving near-lossless compression on coding tasks with Qwen3-Coder-480B and Kimi-K2. The successes of REAP highlight the crucial importance of preserving coordination between the router and experts. Compression methods which impair the router’s ability to independently modulate expert outputs or distort the original functional manifold are less likely to succeed.
@@ 7 Conclusion
Our analysis of current SMoE expert merging techniques introduces irreducible error due to the loss of the router’s independent control over experts. In contrast, expert pruning produces a coordinate subspace of the original layer which maintains the topology of the functional manifold. We introduce REAP, a novel expert pruning method which prunes experts that contribute the least to the layer’s output, thereby minimizing the reconstruction error bound. Empirically, we demonstrate that REAP retains remarkably high accuracy on a wide array of generative tasks across a diverse set of model architectures. We hope that this work inspires further compression techniques for SMoEs and facilitates the deployment of accurate, domain-specific models in resource constrained settings.
@@ Acknowledgments
We would like to acknowledge the helpful feedback of Valavan Manohararajah, Mohammed Adnan, and Rohan Jain. This work was supported in part by RBC Borealis through the RBC Borealis AI Global Fellowship Award. ML and YI gratefully acknowledge the support of Alberta Innovates (ALLRP 577350-22, ALLRP 600038-24), the Natural Sciences and Engineering Research Council of Canada (NSERC) (RGPIN-2022-03120, DGECR-2022-00358), Defence Research and Development Canada (DGDND-2022-03120), and NSERC/Agence Nationale de la Recherche (ANR) (ALLRP 602719-24). This project was supported thanks to funding from IVADO and the Canada First Research Excellence Fund. This research was enabled in part by support provided by the Digital Research Alliance of Canada (alliancecan.ca). YI is supported by a Schulich Research Chair.
@@
This research focused on the algorithmic compression of SMoE models and does not involve the use of human subjects, personally identifiable information, or sensitive data. The datasets used for calibration and evaluation (e.g., C4, evol-codealpaca) are publicly available. Our aim is to enable the use of large-scale SMoE models in resource constrained settings. However, we acknowledge that compression techniques such as REAP could potentially facilitate deployment of models for malicious purposes. Further, our compression methods are applied to pre-trained models and any biases related to fairness, discrimination, or representation inherent in the original models may be present in their compressed versions. We make no attempt in this work to mitigate these potential biases. The primary contribution of this paper is technical, and we do not foresee any new, direct ethical concerns arising from our proposed methodology beyond those already associated with the deployment of large language models.
@@ Appendix A Extension to Hierarchical Clustering
While Equation˜5 analyses pairwise merging, practical implementations often employ hierarchical clustering to form groups of experts. Consider a cluster $C=\{f_{i_{1}},\ldots,f_{i_{k}}\}$ of $k$ experts merged into a single representative $\tilde{f}_{C}$. The original contribution of this cluster can be decomposed as:
@@
After hierarchical merging, the router must apply the summed gate $\sum_{j\in C}g_{i_{j}}$ to a single, static cluster representative $\tilde{f}_{C}$, typically computed as a weighted average of the cluster members based on calibration data. This induces an irreducible error.
Hierarchical clustering error.
@@ Appendix B Additional empirical evidence for loss of independent control
B.1 Functional subspace PCA analysis
Qualitative evidence of functional subspace collapse.
@@ B.1 Functional subspace PCA analysis
In Figure˜1(a), Qwen3’s layer 0 exemplifies the contraction of the functional output space by merging in early layers. The original 128 experts span from 0.4-0.4 to $1.0$ along PC1, pruning maintains this full range with 64 experts, while merging contracts the distribution to approximately [0.2,0.3][-0.2,0.3], a 5-fold reduction. This contraction is dramatic in late layers, where experts are more specialized, as can be seen in Figure˜1(b). Figures˜5(a) and 5(b) exhibit similar contractions of the expert output manifold under merging, whereas pruning often preserves outlier experts and the span of the original expert output manifold.
In Table˜A4, we tabulate the total cumulative variance explained by PC1+PC2 for the PCA projections in Figures˜1 and A5. For low-granularity SMoEs such as Llama-4-Scout and Mixtral, PC1 and PC2 capture most of the variance in the activations. Even in high-granularity SMoEs such as Qwen3 and ERNIE, a large portion of the total variance is captured by PC1 and PC2. The merged variance explained is consistently higher than the baseline, suggesting that the merged outputs have lost some of their high-dimensional complexity. In contrast, the pruned variance explained is consistently lower than the baseline, suggesting that pruning preserves outlier experts and the high-dimensional complexity of the baseline model.
The role of expert granularity.
Both Qwen3-30B-A3B and ERNIE-4.5-21B are highly-granular SMoEs containing 128 and 64 routed experts per layer, respectively, and 8 and 6 routed experts per token, respectively. Functional subspace collapse due to expert merging is more pronounced in these models than in low-granularity models such as Mixtral-8x7B and Llama-4-Scout. With fewer experts and a lower amount of experts per token, low-granularity SMoEs appear to better preserve the variance of their expert output manifolds under expert merging. For example, in Figure˜5(c), the merged manifold spans along PC1 from approximately [0.1,0.2][-0.1,0.2] whereas the pruned manifold spans from approximately $[0.0,0.2]$ along PC1. Similarly, as depicted in Figure˜5(e), the pruned and merged manifolds span PC1 along [0.04,0.0][-0.04,0.0] and [0.05,0.05][-0.05,0.05], respectively.
However, the merged manifold is distorted by the introduction of novel expert functions. For example, in Figure˜5(e), expert merging introduces novel functions which occupy approximately $[0.05,0.005]$ and [0.025,0.005][-0.025,0.005] which are significantly different than any of the original experts. This is best exemplified by Figure˜2, which plots the Wasserstein distance between the original and compressed expert output manifolds in terms of normalized angular distance. Compared to the pruned models, the higher distances between the merged and original expert output manifolds suggest a lower degree of similarity. The distorted manifold of the merged expert outputs represents a loss of fidelity with the original manifold which cannot be restored in the one-shot compression setting.
@@
(e) Llama-4-Scout Layer 0
(f) Llama-4-Scout Layer 47
Figure A5: (5(a),5(c),5(e)) Functional subspace (PCA) for early SMoE layers. Pruning (blue) preserves the manifold geometry; merging (green) collapses it toward the centre. (5(b),5(d),5(f)) Functional subspace (PCA) for late MoE layers.
Table A4: Cumulative variance explained by PC1 and PC2 across compression methods. Compared to pruning, merging results in a consistently higher explained variance suggesting that the merged models have lost some of their high-dimensional complexity.
Model Layer Baseline Merged Pruned
Qwen3-30B-A3B 0 0.2343 0.2700 0.1845
Qwen3-30B-A3B 47 0.7195 0.7437 0.6860
ERNIE-4.5-21B 0 0.3836 0.2851 0.2733
ERNIE-4.5-21B 26 0.2563 0.4599 0.0785
Llama-4-Scout 0 0.9032 0.9343 0.8480
Llama-4-Scout 47 0.9473 0.9546 0.8754
Mixtral-8x7B 0 0.6486 0.8479 0.4016
Mixtral-8x7B 31 0.8580 0.8140 0.7027
B.2 Wasserstein distance
To quantify the distortion of the original expert output manifold, we measure the $1$-Wasserstein distance (Earth Mover’s distance) between the original and compressed expert output manifolds, see Figure˜2.
[…]
where $\mu$ and $\nu$ are uniform probability measures over the indices of $\mathcal{F}$ and $\mathcal$$\hat${F}} respectively, $\Gamma(\mu,\nu)$ is the set of all transport plans (joint distributions) with marginals $\mu$ and $\nu$, and the cost function is defined as the normalized angular distance. This metric quantifies the minimum "work" required to transport the probability mass from the compressed functional manifold to cover the original manifold, thereby penalizing both the contraction of variance (subspace collapse) and the introduction of functionally distinct artifacts (distortion).
@@ Appendix C Evaluation details
Mathematical reasoning is assessed on GSM8K (Cobbe et al., 2021) and MATH-500 (Hendrycks et al., 2021b; Lightman et al., 2023) benchmarks using the evalscope (ModelScope Team, 2024) framework. To assess creative writing, we use 146 creative writing prompts sampled from WildBench (Lin et al., 2024a) with gpt-4o-2024-05-13 used as the judge to evaluate the model responses. We report normalized scores using the WildBench rubric.
@@ Appendix D Baseline methods
The clusters are merged with Equation˜15.
Appendix E Quantization and expert pruning
Pruning is easily combined with quantization, as demonstrated by our large-scale evaluations. We calibrate, prune, and evaluate Qwen3-Coder-480B and Kimi-K2 using FP8 and W4A16 quantization, respectively. In contrast, combining expert merging with quantization is more complex; the fine-grained shared quantization scales across parameter groups must be reconciled during merging, whereas pruning requires no such adjustment
Generally, weight quantization offers superior accuracy at a given compression ratio down to 4-bits. For instance, quantizing Qwen3-30B-A3B to 4-bits using AWQ (Lin et al., 2024b) outperforms a 50% expert-pruned REAP model at 16-bits, despite having half the checkpoint size. However, combining these methods enables compression rates that neither can achieve in isolation. For example, our Kimi-K2 results combine 4-bit weights with 50% expert pruning, representing a total size reduction of 87.5%. Achieving a similar footprint through quantization alone would require 2-bit weights, which are not natively supported on most current accelerators and typically suffer from significant accuracy degradation (Liu et al., 2025b).
To further illustrate this, we calibrate and quantize Qwen3-30B-A3B using the evol-codealpaca-v1 dataset to 4- and 2-bit weights with the LLM Compressor library (AI and vLLM Project, 2024). We additionally compress the 4-bit checkpoint by using REAP calibrated on the same dataset to prune half the experts. In Table˜A5, we report EvalPlus accuracy versus checkpoint size relative to the uncompressed 16-bit model. This preliminary analysis suggests that REAP combined with 4-bit weights outperforms more aggressive quantization. A more thorough exploration of various low-bit schemes and pruning combinations remains an important direction for future work.
Quantization Num. Experts Approx. Relative Size Eval+
W16A16 128 100.0% 81.4
W16A16 64 50.0% 78.0
W4A16-G128 128 25.0% 80.5
W4A16-G128 64 12.5% 77.6
W2A16-G128 64 12.5% 28.6
Table A5: Qwen3-30B-A3B EvalPlus accuracy vs. relative checkpoint size with AWQ quantization, REAP expert pruning, and quantization combined with expert pruning. Overall, quantization outperforms expert pruning up to 4-bit weights. Beyond 4-bit weights, combining quantization with expert pruning yields significantly better accuracy than more aggressive quantization.
Appendix F Additional results
Table˜A6 shows the full suite of MC question answering benchmarks and the average result across all models and methods. Table˜A7 tabulates code generation accuracy of compressed SMoE models calibrated on evol-codealpaca. Eval+ is the average of MBPP+ and HE+. The Code Avg column is the average of Eval+ and LiveCodeBench (LiveCode). Table˜A8 summarizes the accuracy of the various compression methods studied when calibrated with the C4 dataset on coding and MC benchmarks. Notably, while the MC performance is generally slightly higher than models calibrated on evol-codealpaca, the resulting code generation quality is abysmal, with most models failing to generate coherent output.
@@ Appendix F Additional results
REAP 0.354 $\pm$ 0.006 0.503 $\pm$ 0.008 0.737 $\pm$ 0.009 0.481 $\pm$ 0.004 0.496 $\pm$ 0.003 0.309 $\pm$ 0.001 0.561 $\pm$ 0.020 0.584 $\pm$ 0.004 0.503 $\pm$ 0.002
Mixtral-8x7B- Instruct-v0.1 Baseline 0.650 0.842 0.887 0.861 0.691 0.496 0.722 0.740 0.736
@@
REAP 0.546 $\pm$ 0.006 0.782 $\pm$ 0.003 0.841 $\pm$ 0.002 0.777 $\pm$ 0.001 0.565 $\pm$ 0.000 0.458 $\pm$ 0.003 0.739 $\pm$ 0.008 0.677 $\pm$ 0.005 0.673 $\pm$ 0.001
Llama-4-Scout- 17B-16E- Instruct Baseline 0.627 0.848 0.879 0.823 0.803 0.462 0.765 0.692 0.738
@@
REAP 0.418 0.586 0.684 0.628 0.550 0.312 0.581 0.673 0.554
Qwen3-Coder- 480B-A35B- Instruct-FP8 Baseline 0.644 0.822 0.906 0.841 0.850 0.468 0.751 0.717 0.750
@@
Kimi-K2- Instruct- W4A16 Baseline 0.712 0.879 0.913 0.765 0.872 0.504 0.783 0.811 0.780
@@
Table A7: Detailed benchmark results for non-agentic code generation tasks. Eval+ is the average of MBPP+ and HE+. The Code Avg column is the average of Eval+ and LiveCodeBench (LiveCode).